Suppose the first-order maximum for 475 nm blue light being transmitted through a diffraction grating is at an angle of 22°.

Randomized Variables
λ = 475 nm
θ = 22°


How many lines per centimeter are there on this diffraction grating?
1
Expert's answer
2020-12-03T06:59:01-0500

Answer


Equation of m th order maximum

"d\sin \theta =m\lambda"


"d =\frac{m\lambda}{\sin \theta}=\frac{1\times475\times10^{-9}}{\sin 22^°}"


"d=1268\times10^{-9}m"


lines per centimeter is inverse of spacing


"N=\frac{1}{1268\times10^{-9}m}=7886"